# From Intuition to Expertise: Rubric-Based Cognitive Calibration for Human Detection of LLM-Generated Korean Text

### v1

@@ 3.3 Experimental Procedure
We enforce strict information control throughout the three-phase calibration process to preserve experimental integrity. At no point do we disclose the proportion of human-authored versus LLM-generated essays, thereby preventing statistical inference or prior-driven guessing. Each phase employs entirely distinct text samples to eliminate memory-based recognition and ensure that performance gains reflect genuine cognitive calibration rather than familiarity effects.
The transition from Phase 1 to Phase 3 is designed as a cognitive progression: annotators move from intuitive judgments to rubric-guided analysis, and ultimately to domain-specific expertise. This design enables us to test how rubric scaffolding and iterative feedback convert general linguistic competence into specialized detection skill.
By integrating longitudinal exposure, controlled feedback, and analytical justification, our framework operationalizes cognitive calibration as a measurable and reproducible process. This design supports fine-grained analysis of how human experts move beyond superficial fluency to attend to stable, interpretable signals of machine generation.
4 Results of Human Calibration
@@ 4.1 Accuracy Gains Across Calibration Phases
Figure 1 tracks how detection performance evolves under the three-phase calibration protocol. In Phase 1, annotators rely on intuition alone and the human majority vote reaches 60.00% accuracy, which trails the LLM-detector majority vote (96.67%) by 36.67%p. This gap indicates that fluent LLM outputs reliably exploit surface-level plausibility cues that human annotators typically trust when they lack structured decision criteria. Phase 2 introduces rubric-guided judgment and shifts detection from impression-based screening to explicitly justified evaluation. The result shows that this intervention substantially improves accuracy for each annotator and for the majority vote (60% $\rightarrow$ 90%). Appendix B further characterizes Phase 2 with generator-wise score distributions and annotator-level descriptive statistics of total LREAD scores. Phase 3 completes the calibration loop and yields perfect majority-vote performance.
Figure 1: Detection accuracy across the three phases for each human annotator and their majority vote.
@@ 4.2 Error Dynamics and Calibration
Figure 2 shows how calibration reshapes majority-vote errors. In Phase 1, the majority vote yields no false positives (Human mislabeled as AI) but many false negatives (AI mislabeled as Human): the confusion matrix shows 6 true-human essays correctly classified as Human, 0 true-human essays mislabeled as AI, 12 AI essays mislabeled as Human, and 12 AI essays correctly labeled as AI. This pattern suggests that surface-level fluency systematically cues human authorship, causing annotators to mislabel many highly fluent LLM-generated essays as human-written.
@@
Figure 2: Confusion matrices for the human-annotator majority vote in Phase 1 (intuition-only) and Phase 2 (rubric-guided). Each cell reports the number of essays in the corresponding true-label and predicted-label pair under the binary Human versus AI decision.
@@
Figure 3 complements the confusion-based error analysis by visualizing Phase 2 rubric score distributions across generators. By aggregating criteria into Content, Organization, and Expression, the plot frames LREAD as a structured, multidimensional competency profile and reveals systematic source-level variation across dimensions and models.
@@ 4.3 Inter-annotator Agreement Analysis
Table 4 reports Fleiss’ $\kappa$ across phases using the inclusive definition that maps any non-binary response to a third category. We use this formulation because the protocol allows an explicit uncertainty response. This choice aligns with our goal of measuring calibration as a cognitive process: uncertainty is an observable signal in early phases, and the inclusive mapping retains it rather than discarding the corresponding items.
@@
The agreement trajectory mirrors the accuracy gains. Phase 1 yields negative agreement ($\kappa=-0.09$), which indicates systematic inconsistency under intuition-only detection. Phase 2 increases agreement to $\kappa=0.24$, reflecting that rubric-guided criteria align annotators toward shared decision boundaries. Phase 3 reaches strong agreement ($\kappa=0.82$) on the calibrated test set, consistent with a converged expert representation that supports both high accuracy and stable consensus.
Phase N Fleiss’ $\kappa$
Phase 1 (Intuition-only) 30 -0.09
Phase 2 (Rubric-guided) 30 0.24
Phase 3 (Calibrated) 10 0.82
Table 4: Inter-annotator agreement (Fleiss’ $\kappa$) across phases (N = essays per phase). We report the inclusive variant that maps any non-binary response to a third category, preserving a consistent item set while treating abstention or uncertainty as an explicit label.
5 Machine Perception vs. Human Insight
We compare the calibrated human signatures with signatures produced by three LLM-based reviewers after Phase 1. This contrast isolates which cues are salient to strong language models under zero-shot inspection and which cues become diagnostic through human calibration. We provide the consolidated, model-wise human signature inventory in Appendix C and focus here on the comparative analysis with LLM detectors.
@@ 5.1 Human–LLM Comparison Protocol
The comparison contrasts two qualitatively different sources of evidence. The human signatures were elicited after Phase 3, after annotators received gold labels for all phases and completed rubric-based calibration. The LLM signatures were produced after Phase 1, before rubric construction, reflecting zero-shot inspection by strong language models. This setup helps disentangle broadly accessible cues from calibration-specific micro-diagnostics that emerge through iterative feedback.
Cue Family Human Annotators LLM Detectors Relation
Macro-structure Model-specific templates (e.g., enumeration and conclusion openings); template mixing becomes diagnostic for some models. Emphasis on canonical 3-part structure and balanced concession patterns; “too complete” structure as an AI cue. Overlap
Surface mechanics Optional-but-recurrent comma placement around connectives and clause endings; spacing irregularities; within-essay inconsistency as a cue. Mechanics used mainly as a coarse cue: near-zero typos implies AI; disfluency clusters imply human. Divergence
Lexicon and register Sino-Korean density and nominalization as discriminators; two types of translationese identified—Sino-Korean calques (Qwen2) versus English-style modifier-noun constructions (GPT-4o). High-level register labels (report-like tone, mature vocabulary) without pinpointing micro-realization patterns. Partial
Persona and affect Writer persona becomes diagnostic; collective responsibility and forward-looking closures recur as AI cues. Emotion and lived experience as human cues; safe neutrality and socially agreeable resolutions as AI cues. Overlap
Evidence style Concrete elaboration and counterargument handling differentiate models; generic remedies appear suspicious. Personal anecdotes and messy constraints imply human; generic institutional prescriptions imply AI. Overlap
Table 5: Comparison of cues emphasized by calibrated human annotators and LLM detectors. “Relation” indicates whether the cue family primarily overlaps or diverges in operational granularity.
5.2 Translationese: A Human Detection Signal
As summarized in Table 5, the contrast between calibrated human annotators and LLM-based reviewers reveals both substantive overlap and systematic divergence across cue families. At a high level, both groups attend to macro-structural regularities and generic resolution patterns, suggesting that certain discourse-level signals of machine authorship are broadly accessible even under zero-shot inspection. However, the table also highlights a consistent asymmetry in operational granularity: calibrated humans attend to stylistically optional but model-specific punctuation habits, such as the excessive use of commas after sentence-initial connectives or clause-linking endings, which LLM detectors frequently overlook as surface fluency. This divergence motivates the subsequent analysis, which distinguishes shared cues that emerge without calibration from translationese-like and punctuation-level phenomena that become discriminative only through iterative human feedback.
5.3 Shared Cues
Humans and LLM evaluators broadly converged on macro-regularity as a primary signal of machine authorship. Both groups identified a highly standardized “General-to-Specific” logic flow and a mechanical 3-part structure, particularly in GPT-4o, which consistently utilizes ordinal markers (e.g., “First”, “Second”) to partition arguments. Both also treated repetition—either at the level of connectives (Solar) or at the level of sentence-final realizations—as a strong cue, and both observed that LLM essays often terminate in broadly agreeable prescriptions (e.g., education or institutional reform) rather than topic-specific, lived-experience constraints.
5.4 Divergent Cues
The primary divergence concerns granularity. Following calibration, human annotators relied on Korean-specific micro-patterns that resist reduction to generic fluency judgments, including optional comma placement after sentence-initial connectives, punctuation following clause-linking endings, spacing at multiword boundaries, and register diagnostics associated with Sino-Korean nominalization. In contrast, LLM detectors emphasized global properties such as overall coherence, balanced dialectical framing, and the safety of examples and solutions. Although these properties often reflect shared underlying regularities, they operate at a coarser granularity and do not explicitly encode language-specific mechanisms.
Human annotators consistently observed a translated quality in model outputs, with Qwen2 displaying Sino-Korean nominalization and calque-like phrasing, and GPT-4o exhibiting English-style translationese characterized by modifier-heavy noun phrases and frequent pronominal reference. Both patterns were grammatically well-formed yet rhetorically non-native. Prior work has shown that translationese carries systematic source-language fingerprints and stylistic shifts that can make translated text distributionally distinct from natively authored text (doshi-etal-2024-pretraining). Evaluation pipelines that rely on translated instructions can further confound naturalness, complicating fair comparisons between English-conditioned and target-language behaviors (enomoto-etal-2025-fair).
Recent studies connect these artifacts to multilingual LLM training and alignment. li-etal-2025-lost show that supervised training can shape literal, translation-like realizations, and guo-etal-2025-large report an “English accent” effect in multilingual generation that degrades perceived target-language naturalness. We therefore treat translationese-like interference as a discriminative cue family that calibrated humans can exploit, and we provide an expanded discussion with concrete measurement candidates in Appendix D.
6 Related Work
6.1 Post-hoc Detection & LLM Watermarking
Prior work on attributing LLM-generated text has largely followed two directions: post-hoc detection (mitchell2023detectgpt; su-etal-2023-detectllm; bao2024fastdetectgpt; zhu-etal-2023-beat; yang2024dnagpt) and LLM watermarking (kirchenbauer2023watermark; dathathri2024scalable; lu-etal-2024-entropy; chen-etal-2024-watme; hou-etal-2024-semstamp; guo2024context; wang-etal-2025-morphmark; park2026watermod).
Post-hoc detection methods typically operate without altering the source model, making them suitable for retrospective auditing. DetectGPT (mitchell2023detectgpt) exploits probability curvature in the likelihood landscape, with subsequent work improving efficiency via log-rank–based estimators (su-etal-2023-detectllm) and conditional probability curvature (bao2024fastdetectgpt). Other approaches leverage LLM querying through revision-based similarity analysis (zhu-etal-2023-beat) or analyze divergent continuations and n-gram distributions (yang2024dnagpt). Despite strong empirical performance, these methods remain sensitive to domain shift and continued advances in generation quality.
LLM watermarking approaches address these limitations by embedding attribution signals directly into the decoding process. kirchenbauer2023watermark introduce a foundational logit-bias watermarking framework, which subsequent work extends toward scalable deployment (dathathri2024scalable) and a range of design variants, including entropy-aware (lu-etal-2024-entropy), lexical (chen-etal-2024-watme), semantic (hou-etal-2024-semstamp), and contextual (guo2024context) schemes. MorphMark (wang-etal-2025-morphmark) advances this line by adaptively modulating watermark strength using the cumulative probability mass of green-list tokens to address the effectiveness–quality trade-off. WaterMod (park2026watermod) refines the paradigm through probability-ranked modular partitioning. Collectively, these approaches establish strong baselines while treating attribution as automated inference, with limited emphasis on human judgment.
6.2 Linguistic Profiling of AI-Generated Text
zanotto-aroyehun-2025-linguistic find that LLM-generated text often shows deeper and longer dependency structures than human writing, while human texts exhibit more diverse semantic content and greater overall variation. They further note that newer models become more human-like on syntactic depth and length metrics, shrinking earlier gaps. Within the Korean context, park-etal-2025-katfishnet identify language-specific artifacts in spacing, morphology, and punctuation that differentiate LLM-generated Korean text from human writing. These observations highlight the importance of attending to fine-grained linguistic realizations beyond surface-level fluency.
6.3 Human Expertise in AI Detection
russell-etal-2025-people demonstrate that frequent LLM users outperform many automatic detectors when identifying synthetic English articles, while providing interpretable rationales grounded in linguistic cues. dugan2023real demonstrate that boundary detection remains challenging and exhibits substantial inter-annotator variability; nevertheless, targeted interventions are shown to produce measurable improvements over time.
@@ 7 Conclusion
We demonstrate that human attribution of LLM-generated text is a learnable and calibratable competency, not a fixed bottleneck of expert judgment. In a three-phase longitudinal study, linguistically trained annotators transition from surface-fluency heuristics to LREAD-scaffolded analysis, achieving substantial accuracy gains with markedly higher inter-annotator agreement and improved sensitivity to language-specific signals that automated detectors still struggle to capture reliably.
@@
The empirical evaluation focuses exclusively on argumentative essays at the primary and secondary school levels, a genre that foregrounds discourse organization and reasoning coherence. This controlled setting enables precise measurement of how rubric scaffolding reshapes judgments beyond surface fluency, but it leaves open whether the same calibration trajectory holds for genres with different stylistic conventions—conversational dialogue, narrative fiction, or technical documentation. Extending LREAD to these domains will require principled rubric adaptation and systematic revalidation.
@@
Appendix A LREAD Rubric Specifications
@@
Appendix B Phase 2 Rubric Score Distributions and Annotator-Level Descriptives
This appendix complements the main Phase 2 results by (i) visualizing generator-wise distributions of total rubric scores and (ii) reporting annotator-level descriptive statistics that contextualize grade-level and generator-level variability under the same scoring protocol.
B.1 Total LREAD Score Distributions

### v2

@@ 4.3 Experimental Procedure
We maintain strict information control by withholding label distributions and using disjoint essay sets across phases. This design limits memorization and distributional inference, helping isolate calibration effects from familiarity.
Annotators progress from intuition-only screening to rubric-guided analysis and then to internalized application on a small held-out subset. This phased design allows us to assess how explicit criteria affect individual attribution decisions and panel-level consistency.
5 Calibration Outcomes
@@ 5.1 Accuracy Gains Across Calibration Phases
Figure 1 summarizes majority-vote performance across the three phases. In Phase 1, intuition-only attribution reaches 0.60 accuracy on 30 essays and is not significantly above chance, illustrating the severity of fluency-driven error before calibration. In Phase 2, rubric-guided attribution improves to 0.90 on a disjoint 30-essay set, indicating that criterion-anchored evaluation substantially improves performance. In Phase 3, the panel classifies all 10 essays correctly on a limited held-out elementary-persona subset; however, the exact 95% confidence interval [0.692, 1.000] indicates substantial residual uncertainty. We therefore interpret Phase 3 as encouraging small-set evidence of successful internalized application rather than a population-level estimate of perfect human detection.
[…]
Figure 1: Detection accuracy across the three phases for each human annotator and for the majority vote. Exact 95% confidence intervals are shown for the majority vote (Phase 1: $N=30$; Phase 2: $N=30$; Phase 3: $N=10$). The black dashed line denotes the majority-vote performance of the LLM detectors in Phase 1.
@@ 5.2 Error Dynamics and Calibration
Figure 2: Confusion matrices for the human-annotator majority vote in Phase 1 (intuition-only) and Phase 2 (rubric-guided), with $N=30$ essays per phase. Each cell reports the number of essays in the corresponding true-label and predicted-label pair under binary Human-versus-AI attribution.
Figure 2 shows how calibration reshapes majority-vote errors. In Phase 1, majority voting yields no false positives (Human mislabeled as AI) but a considerable number of false negatives (AI mislabeled as Human). The confusion matrix shows that all 6 human-written essays are correctly classified as Human, whereas among AI-generated essays, 12 are misclassified as Human and 12 are correctly classified as AI. This pattern suggests that surface-level fluency systematically cues human authorship, causing annotators to mislabel many highly fluent LLM-generated essays as human-written under intuition-only screening.
@@
This asymmetric shift is important. If calibration merely induced a generalized AI-leaning response bias, we would expect a much larger increase in false positives. Instead, false positives remain nearly unchanged, whereas false negatives on AI essays drop sharply from 12 to 2. This pattern is more consistent with a targeted correction of the fluency-trap failure mode than with blanket over-detection. The reduction in false negatives is statistically significant under Fisher’s exact test ($p=0.003$).
@@
Figure 3 complements the confusion-based error analysis by visualizing Phase 2 rubric score distributions across generators. By aggregating criteria into Content, Organization, and Expression, the plot frames LREAD as a structured, multidimensional competency profile and reveals systematic source-level variation across dimensions and models. In particular, the figure shows that human essays remain comparatively weaker on formal expression-related dimensions, whereas several LLM sources cluster at higher levels of rubric conformity, reinforcing our central claim that surface correctness can be a misleading signal of human authorship.
@@ 5.3 Inter-Annotator Agreement Analysis
Phase NN Accuracy 95% CI (exact) TN / FP / FN / TP FN rate Fleiss’ $\kappa$ Unanimous
Phase 1 30 0.600 [0.406, 0.773] 6 / 0 / 12 / 12 50.0% (12/24) 0.088-0.088 5/30
Phase 2 30 0.900 [0.735, 0.979] 5 / 1 / 2 / 22 8.3% (2/24) 0.243 18/30
Phase 3 10 1.000 [0.692, 1.000] 2 / 0 / 0 / 8 0.0% (0/8) 0.820 9/10
Table 4: Phase-wise summary of majority-vote performance and inter-annotator agreement. We report accuracy with exact 95% confidence intervals, majority-vote confusion counts (TN / FP / FN / TP), the false-negative rate on AI essays, inclusive Fleiss’ $\kappa$, and the number of unanimously labeled essays. Fleiss’ $\kappa$ uses the uncertainty-aware inclusive definition that maps any non-binary response to a third category.
@@
Table 4 summarizes phase-wise majority-vote performance together with inter-annotator agreement. We report Fleiss’ $\kappa$ using the inclusive definition that maps any non-binary response to a third category, because our protocol allows an explicit uncertainty response. This formulation is appropriate for the present setting: uncertainty is itself an observable part of the calibration process, especially in earlier phases, and the inclusive mapping preserves these responses rather than removing them from the agreement analysis.
@@
The agreement trajectory broadly mirrors the gains in attribution performance. Phase 1 yields negative agreement ($\kappa=-0.088$), indicating systematic inconsistency under intuition-only detection. Phase 2 increases agreement to $\kappa=0.243$, suggesting that rubric-guided criteria align annotators toward more stable and shared decision boundaries. Phase 3 reaches strong agreement ($\kappa=0.820$) on the held-out calibrated set, together with near-complete unanimous labeling (9/10), consistent with a consolidated evaluative framework.
Fleiss’ $\kappa$ alone, however, does not explain why disagreement occurs. A more informative pattern is found in directional ambiguity: in Phase 1, annotators often appeared to identify similar cue families, yet assigned them to opposing labels, resulting in systematic disagreement rather than random noise. After calibration, disagreement becomes both less frequent and less directionally unstable, which suggests that the rubric helps align not only what annotators attend to, but also how they interpret those cues in attribution decisions.
We therefore interpret the rise in agreement as evidence of criterion alignment rather than simple convergence through repetition. This interpretation is also consistent with the table’s broader phase-wise pattern: as calibration proceeds, false negatives on AI essays drop sharply, overall accuracy improves, and unanimous decisions become substantially more common. Appendix B reports pairwise agreement tables, phase-level uncertainty summaries, and inferential tests.
6 Auditable Human Calibration vs. Zero-Shot LLM Reviewing
This comparison identifies which attribution cues remain inspectable and maintainable under human calibration, relative to zero-shot LLM review. Although strong LLM judges can perform well under blind prompting, their decisions are often opaque and vulnerable to provider-side drift. In contrast, LREAD makes attribution reasoning explicit, reviewable, and auditable.
We separate generator and judge models to avoid self-recognition confounds. Generators define the essay sources, and judge models provide an external zero-shot review baseline, better reflecting deployment with heterogeneous generators.
Dimension Calibrated human cues Zero-shot LLM-reviewer cues Interpretation
Content generic institutional solutions; weak personal constraint; safe argumentation; persona inconsistency generic neutrality; socially agreeable resolutions; emotionally flat closure Overlap at a macro level; humans are more persona-sensitive
Organization over-regularized scaffold; too-complete paragraph structure; standardized conclusion pivots canonical three-part structure; balanced concession patterns; “too complete” organization Strong overlap
Expression spacing regularity; optional comma habits; register stability; natural-Korean-phrasing issues; translationese micro-patterns near-zero typos; mature/report-like tone; broad fluency cues Strongest divergence in granularity
Table 5: Comparison of cues emphasized by calibrated human annotators and zero-shot LLM reviewers.
@@ 6.1 Human–LLM Comparison Protocol
This comparison brings together two different sources of evidence. Human signatures are elicited after Phase 3, once annotators complete calibration and receive gold labels. LLM signatures, by contrast, come from zero-shot reviewing elicited only after an initial blind O/X decision. We therefore interpret this analysis as a contrast between calibration-enriched human diagnostics and broadly accessible zero-shot LLM cues, rather than as a fairness-controlled contest. Exact blind-attribution prompts, post-hoc cue-elicitation prompts, and voting conventions are reported in Appendix C.
6.2 Translationese and Operational Granularity
As summarized in Table 5, human annotators and LLM reviewers overlap in macro-structural cues while diverging in operational granularity. Both groups attend to regular discourse structure and generic resolution patterns, suggesting that some signals of machine authorship remain accessible under zero-shot inspection.
The key difference is that calibrated humans treat punctuation, spacing, register stability, and natural Korean phrasing distortions as explicit diagnostic evidence, whereas LLM reviewers often subsume these patterns under broad fluency judgments. Such cues matter in practice only when they can be taught, inspected, and revised over time. The full after-calibration signature inventory is reported in Appendix D, including model-specific form-level and content-level cues.
6.3 Shared and Divergent Cues
Humans and zero-shot LLM evaluators converged on macro-regularity as a signal of machine authorship. Both groups identified standardized logic flow, mechanically balanced three-part organization, repetition, and generic agreeable conclusions as cues that a text had been generated rather than written under topic-specific or persona-specific constraints.
The main divergence concerned granularity. After calibration, human annotators relied on Korean-specific micro-patterns, including optional comma placement, punctuation after clause-linking endings, spacing at multiword boundaries, and register diagnostics associated with Sino-Korean nominalization. By contrast, LLM detectors emphasized broader properties such as coherence, balanced framing, and the generic safety of examples and solutions. Model-specific patterns sharpened this contrast: Qwen2 was associated with Sino-Korean nominalization and calque-like phrasing, whereas GPT-4o more often exhibited English-style translationese, including modifier-heavy noun phrases and frequent pronominal reference. Prior work links such artifacts to multilingual LLM training and reduced target-language naturalness (Doshi et al., 2024; Enomoto et al., 2025; Li et al., 2025; Guo et al., 2025). We therefore treat translationese-like interference as a discriminative cue family exploited by calibrated humans, and Appendix E outlines how this qualitative cue family could be operationalized in future work.
@@ 7 Conclusion
We provide pilot evidence that rubric-based calibration can improve human attribution of LLM-generated Korean argumentative essays under tightly controlled conditions. Across our three-phase study, linguistically trained annotators move from intuition-only screening to criterion-anchored analysis, with gains in accuracy, agreement, and reasoning transparency. More broadly, LREAD contributes a way of making attribution judgments explicit and auditable by turning diffuse intuitions into an evaluative practice that can be taught, inspected, and revised. In this sense, the value of the framework lies not only in improved performance within our controlled study, but also in offering a human-centered complement to automated detectors in settings where transparency, auditability, and maintainability matter.
@@
The empirical evaluation focuses exclusively on argumentative essays at the primary and secondary school levels, a genre that foregrounds discourse organization and reasoning coherence. Although this controlled setting enables a precise assessment of how rubric-based scaffolding reshapes judgments beyond surface-level fluency, it does not establish whether the same calibration trajectory generalizes to genres with distinct stylistic conventions, including conversational dialogue, narrative fiction, and technical documentation. Extending LREAD to these domains will require principled rubric adaptation and systematic revalidation.
@@
Appendix A Full LREAD Rubric and Detection-Oriented Adaptations
This section documents the full LREAD rubric as used in this study. We distinguish between (i) stable, framework-level dimensions that are likely to transfer across settings (Content, Organization, Expression) and (ii) adaptable micro-diagnostics that are intentionally sensitive to language, genre, and writer-persona conditions (e.g., orthography, spacing, register, and natural Korean phrasing / translationese-sensitive judgments). This distinction is central to interpreting LREAD as a human-centered calibration methodology instantiated in Korean rather than as a fixed universal rubric. The framework-level structure is designed to remain conceptually stable, whereas the micro-diagnostics are expected to be refreshed and revalidated when the target language, genre, or generator population changes.
Relative to the source writing standard, LREAD introduces four key adaptations for attribution: (i) finer expression-level granularity, (ii) explicit natural-Korean-phrasing judgments, (iii) dual creativity criteria at content and expression levels, and (iv) score anchors tailored to one-paragraph student essays rather than longer compositions. These adaptations clarify how LREAD differs from a conventional writing-assessment rubric and why it functions as an attribution-oriented human calibration instrument rather than as a generic composition-scoring tool.
@@ Appendix A Full LREAD Rubric and Detection-Oriented Adaptations
Appendix B Performance, Statistical, and Rubric Diagnostics
This appendix consolidates quantitative support that is too detailed for the main narrative but essential for interpreting the study responsibly. It reports individual trajectories, uncertainty-aware phase-level performance, pairwise agreement, inferential statistics, and rubric-score descriptives. Together, these materials complement the main text by showing not only that calibration improves panel-level outcomes, but also how those improvements distribute across annotators, error types, and rubric dimensions.
B.1 Individual Annotator Accuracy Trajectories
Table 6 reports phasewise accuracy for each annotator and for the majority vote. These trajectories help separate panel-level improvement from idiosyncratic individual fluctuation.
Annotator Phase 1 ($N=30$) Phase 2 ($N=30$) Phase 3 ($N=10$)
A1 50.0% (15/30) 66.7% (20/30) 90.0% (9/10)
A2 46.7% (14/30) 96.7% (29/30) 100.0% (10/10)
A3 80.0% (24/30) 86.7% (26/30) 100.0% (10/10)
Majority vote 60.0% (18/30) 90.0% (27/30) 100.0% (10/10)
Table 6: Phasewise accuracy trajectories for individual annotators and the majority vote.
B.2 Majority-Vote Performance with Uncertainty Quantification
Table 7 reports phase-level outcomes with exact confidence intervals, confusion counts, agreement estimates, and unanimity counts. This table should not be read as evidence of population-level ceiling performance; rather, it shows that the strongest phase-level gain occurs from Phase 1 to Phase 2, while Phase 3 provides encouraging but small-sample evidence of internalized application.
Phase NN Accuracy 95% CI (exact) TN / FP / FN / TP Fleiss’ $\kappa$ $\kappa$ 95% CI Unanimous
1 30 0.600 [0.406, 0.773] 6 / 0 / 12 / 12 0.088-0.088 [0.254-0.254, 0.075] 5/30
2 30 0.900 [0.735, 0.979] 5 / 1 / 2 / 22 0.243 [0.030, 0.473] 18/30
3 10 1.000 [0.692, 1.000] 2 / 0 / 0 / 8 0.820 [0.071-0.071, 1.000] 9/10
Table 7: Majority-vote performance across phases with uncertainty quantification. Accuracy intervals are exact 95% binomial confidence intervals. Confidence intervals for Fleiss’ $\kappa$ are based on item-level bootstrap resampling (B=20, [^0] B=20{,}000). Because Phase 3 contains only 10 essays, its bootstrap interval for $\kappa$ is unstable and should be interpreted together with the unanimity count.
B.3 Pairwise Agreement and Directional Ambiguity
To complement Fleiss’ $\kappa$, Table 8 reports pairwise Cohen’s $\kappa$ across phases. The most informative pattern is the transition from directional opposition to alignment, especially for the Annotator 2–Annotator 3 pair.
Annotator Pair Phase 1 Phase 2 Phase 3
A1–A2 0.14 0.03-0.03 0.75
A1–A3 0.14 0.03 0.75
A2–A3 0.53-0.53 0.73 1.00
Table 8: Pairwise Cohen’s $\kappa$ across phases. The A2–A3 trajectory is consistent with directional ambiguity in Phase 1 and criterion alignment after calibration.
B.4 Statistical Tests